OpenStudy (anonymous):

How to solve and answer 15-3.7m=12.84-2.5m

jigglypuff314 (jigglypuff314):

you want to get m alone on one side of the equals sign, so to solve for m for 15-3.7m=12.84-2.5m you first add 3.7m to both sides so +3.7m +3.7m 15 - 0 = 12.84 + 1.2m then you subtract both sides by 12.84 -12.84 -12.84 2.16 = 1.2m then divide both sides by 1.2 /1.2 /1.2 m = 2.16/1.2 so m =

OpenStudy (anonymous):

oh okay thank you